7.34 Cyclopropane, C3H6, is an anesthetic given by inhalation. How many grams are in 0.25 mol of cyclopropane?
1
Expert's answer
2018-11-23T03:19:58-0500

Answer on Question #83263, Chemistry/Organic Chemistry

7.34 Cyclopropane, C3H6, is an anesthetic given by inhalation. How many grams are in 0.25 mol of cyclopropane?

Answer

n=m/M

m = M*n

M (C3H6) = 12.01*3+1.01*6 =42.09 (g/mol)

n(C3H6) = 42.09 g/mol *0.25 mol = 10.52 g

Answer: 10.52 g